(a) An electron moves along a circle of radius 1 m in a perpendicular magnetic field of strength 0.50 T. What would be its speed? Is it reasonable? (b) If a proton moves along a circle of the same radius in the same magnetic field, what would be its speed?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. (a) R = 1 n, B= 0.5 T, r = mυ/qB ⇒ 1 = 9.1 * 10^-31 * υ/1.6 * 10^-19 * 0.5 ⇒ υ = 1.6 * 0.5 * 10^-19/9.1 * 10^-31 = 0.0879 * 10^10 = 8.8 * 10^10 m/s No, it is not reasonable as it is more than the speed of light. (b) r = mυ/qB ⇒ 1 = 1.6 * 10^-27 * v/1.6 * 10^-19 * 0.5 ⇒ υ = 1.6 * 10^-19 * 0.5/1.6 * 10^-27 = 0.5 * 10^8 = 5 * 10^7 m/s.